OpenStudy (anonymous):

Which system has (-8, 0) as the solution? a. x - 4y = -8 2x - 3y = -16 b. x - y = -2 3x - y = 0 c. x + y = -4 -3x + 2y = 2 d. x + y = -1 4x + 4y = -4 e. 4x + 6y = -12 2x + 3y =

OpenStudy (astrophysics):

Plug in the values (-8,0) in the equations and see which is true

OpenStudy (astrophysics):

So looking at b for example, x-y=2 -8-0=-2 this is obviously wrong right? So we can eliminate this one.

OpenStudy (anonymous):

yes, thank you i got it.